OpenStudy (anonymous):

find first 15 term of geometric sequence of 4,12,36,108

OpenStudy (ranga):

Can you figure out the common ratio of the geometric series?

OpenStudy (anonymous):

no

OpenStudy (ranga):

This is how a geometric series is formed. You start with the first number. Here it is 4. You multiply the first number by another number called the common ratio to get the second number. You multiply the second number by the common ratio to get the third number. And so on and so forth. So the first number 4 was multiplied by some number (call it r) to get the second number which is 12. Can you tell me what r is?

OpenStudy (anonymous):

3

OpenStudy (ranga):

Yes. The common ratio = 3. First number = 4 Second number = 4 x 3 = 12 Third number = 12 x 3 = 36 Fourth Number = 36 x 3 = 108 Fifth Number = ?

OpenStudy (anonymous):

324?

OpenStudy (ranga):

Yes. Sixth Number = ?

OpenStudy (anonymous):

972

OpenStudy (ranga):

Yes. Keep doing it until you get 15 numbers. The geometric sequence so far are: 4, 12, 36, 108, 324, 972, that covers 6 numbers and you have 9 more to go to get 15 numbers.

OpenStudy (anonymous):

i think is 19,131,876

OpenStudy (anonymous):

so will these be the sum of the 15 term?

OpenStudy (ranga):

No, not the sum. They want you to just list the first 15 terms. Not add them up.

OpenStudy (anonymous):

what if they do?

OpenStudy (anonymous):

oh ok i check they actually do ask for it

OpenStudy (ranga):

There is a standard formula for the sum of the first n terms of a geometric series. You can look it up in a text book or I can give it to you. You simply plug in the values into the formula to get the sum. But here they simply want you to list all the 15 terms of the geometric series. The number that you gave earlier, 19,131,876 is the 15th term of the series. So you have to list all the terms from 4, 12, 36, 108, 324, 972, ...., ...., ...., 19131876. Also don't put commas in 19131876 because that will be confusing as the series has commas to separate each term.

OpenStudy (anonymous):

can u give to me?

OpenStudy (ranga):

Sum = \[a \frac{ (r ^{n} - 1) }{(r - 1) }\] where a is the first term in the geometric sequence; r is the common ratio; n is the number of terms in the geometric sequence. Plug in the values for a, r and n and you will get the sum.

OpenStudy (anonymous):

so is it 36?

OpenStudy (ranga):

You are getting sum = 36?

OpenStudy (anonymous):

or is it the root 4374

OpenStudy (ranga):

No. The sum will be a very large number. Remember you yourself listed the 15th term in the sequence as: 19,131,876. So the sum of the first 15 terms will be even larger than that.

OpenStudy (ranga):

Substitute a = 4, r = 3, n = 15 in the formula for the sum.

OpenStudy (ranga):

I am getting sum = 28,697,814

OpenStudy (anonymous):

ok can u solve this find A 250 when a1 = -50,d=5?

OpenStudy (ranga):

The problem is not typed correctly. Missing information.

OpenStudy (anonymous):

use the formula of general term (the nth term) of arithmetic sequence to find the in dicated term of the sequence with the given first term, a1, and common difference, d so find a250 when a1= -50, d=5

OpenStudy (anonymous):

can u still help?

OpenStudy (ranga):

The nth term of an arithmetic sequence: a(n) = a(1) + (n-1)d a(1) = -50, d = 5, n = 250 a(250) = -50 + (250 - 1)5 = -50 + 1245 = 1195.